OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (anonymous):

why

OpenStudy (gorv):

|x-2|=12-10=2

OpenStudy (igreen):

You have to solve two equations: x - 2 + 10 = 12 and x + 2 + 10 = 12 ---------------------- x - 2 + 10 = 12 Add: x + 8 = 12 Subtract 8 to both sides: x = 4 ----------------------------- x + 2 + 10 = 12 Add: x + 12 = 12 Subtract 12 to both sides: x = 0

OpenStudy (gorv):

now mode will alwys give positive value

OpenStudy (igreen):

So your answer is x = 4, x = 0(Option A).

OpenStudy (gorv):

so to remove mode to condition will be there \[x-2=\pm 2\]

OpenStudy (gorv):

x-2=2 -----> x=2+2=4--->X=4 x-2=-2----->X=-2+2---->X=0

OpenStudy (anonymous):

what is the anwer